i need help with this problem.
-3x+4y=11
rewrite as a function of X
please help.

Answer by tutorcecilia(2144)

-3x+4y=11 [solve for the y-term]
-3x+3x+4y=3x+11
4y=3x+11
4y/4=3x/4+11/4
y=3/4x+11/4
or f(x)=3/4x+11/4